# Francisco Calvillo LPSM, Sorbonne Université, 4 Place Jussieu, 75005 Paris, France Luc Devroye

### v1

@@ 4.1 Moderate tails
Since $Y_{n}\to\infty$ in probability, $I_{n}\to\infty$ in probability as well. Moreover, for all $n\geq 1$, one can choose $N_{n}$ such that $N_{n}-n$ is distributed as $N_{0}$. The fact that all vertices $k\geq N_{n}$ belong to the tree rooted at $I_{n}$ in this single-choice model implies that $I_{n}\in T_{k}$ for all $k\geq N_{n}$, and in particular $L_{I_{n}}\leq L_{k}$ for all $k\geq N_{n}$. From Proposition 11 we know that $L_{n}\to\infty$ in probability. Hence, $L_{I_{n}}\to\infty$ in probability. Moreover, for any fixed $m\geq 1$ we have
$\mathbb{P}\left(\cup_{k\geq 2n}\left[L_{k}\leq m\right]\right)$ $\leq\mathbb{P}\left(N_{n}>2n\right)+\mathbb{P}\left(N_{n}\leq 2n,L_{I_{n}}\leq m\right)$
$\leq\mathbb{P}\left(N_{0}>n\right)+\mathbb{P}\left(L_{I_{n}}\leq m\right)~{}.$
Thus, for any $m\geq 1$, $\mathbb{P}\left(\cup_{k\geq 2n}[L_{k}\leq m]\right)\to 0$ as $n\to\infty$. By continuity of measure, this implies that for any $m\geq 1$,
$\mathbb{P}\left(\cap_{n\geq 1}\cup_{k\geq 2n}\left[L_{k}\leq m\right]\right)=0.$
@@
$\mathbb{P}\left(\cup_{m\geq 1}\cap_{n\geq 1}\cup_{k\geq 2n}\left[L_{k}\leq m\right]\right)=0~{},$
which implies that $L_{n}\to\infty$ almost surely.
@@ 4.2 Heavy tails
In this section we study the heavy-tailed case, that is, $\mathbb{E}\min(Z,W)=\infty$. From Theorem 3 we already know that $L_{n}$ cannot go to infinity with probability one. Here we show that, in contrast to this, $L_{n}\to\infty$ in probability.
@@
$\mathbb{P}\left(-k\in T_{n}^{Z}\right)$ $=\sum_{i=1}^{n}\mathbb{P}\left(Z_{i}=i+k\right)\mathbb{P}\left(i\in T_{n}\right)$
$=\sum_{i=1}^{n}q_{k+i}r_{n-i}$
$=\sum_{i=k+1}^{n+k}q_{i}r_{n+k-i}$
$\leq\sum_{i=1}^{n+k}q_{i}r_{n+k-i}=r_{n+k}~{}.$
Proof.  [Proof of Theorem 7] Fix a constant $w>0$ and let $I_{n}^{w}$ denote the number of intervals of the form $(kw,(k+1)w]$, with $k\in\mathbb{Z}_{\geq 0}$, that are intersected by the chain $T_{n}^{Z}$. Note that, for any $x>0$, if $I_{n}^{w}<x$ then $T_{n}^{Z}<wx$, thus
@@
$\quad\leq\mathbb{E}\sum_{\begin{subarray}{c}k,m\in A_{n}\\ \textit{s.t. }k\neq m\end{subarray}}\mathbbm{1}_{T_{k}^{W}\cap T_{m}^{W}\neq\emptyset}$
$\quad\leq\mathbb{E}\sum_{\begin{subarray}{c}k,m\in A_{n}\\ \textit{s.t. }k\neq m\end{subarray}}\mathbb{P}\left(T_{k}^{W}\cap T_{m}^{W}\neq\emptyset\right)$
@@
$\mathbb{P}\left(T_{k}^{W}\cap T_{m}^{W}\neq\emptyset\right)$ $\leq\sum_{i=-\infty}^{k}\mathbb{P}\left(i\in T_{k}^{W}\right)\mathbb{P}\left(i\in T_{m}^{Z}\right)$
$\leq\sum_{i=-\infty}^{k}r_{k-i}r_{m-i}$
$\leq\sqrt{\sum_{i=0}^{\infty}r_{i}^{2}}\times\sqrt{\sum_{i=m-k}^{\infty}r_{i}^{2}}.$ (4.19)
@@
$\mathbb{P}\left(\cup_{k\neq m,k\in A_{n}}T_{k}^{W}\cap T_{m}^{W}\neq\emptyset\right)\leq{\ell\choose 2}\sqrt{\sum_{i=0}^{\infty}r_{i}^{2}}\times\sqrt{\sum_{i=w}^{\infty}r_{i}^{2}}.$ (4.20)
@@
Finally, observe that if $L_{n}<\ell$ and $I_{n}^{w}\geq 2\ell$, then there have to be at least two distinct vertices $k,m$ in $A_{n}$ such that the chains $T_{k}^{W}$ and $T_{m}^{W}$ meet each other: otherwise, each chain $T_{k}^{W}$ with $k\in A_{n}$ would lead to a distinct leaf in $T_{n}$, contradicting the event $L_{n}<\ell$. Thus, by (4.20) we have
$\mathbb{P}\left(L_{n}<\ell\right)\leq\mathbb{P}\left(I_{n}^{w}<2\ell\right)+\epsilon~{}.$
@@
$\limsup_{n\to\infty}\mathbb{P}\left(L_{n}<\ell\right)\leq\epsilon~{},$
@@ 5 Conclusion
We introduced and studied a simple mathematical model for recommendation systems based on giving each user two random choices, which include a mixture of past choices and untried options. In this setup, we identify three regimes based on the tail behavior of the sizes of the jumps in the past. A natural generalization of the model allows $k>2$ choices, left for further study. We anticipate that in that case there are $k+1$ regimes, but even the case $k=3$ seems interesting. The eagle-eyed reader must have noticed that the bottom right corner of the table summarizing the results is left unfilled. Indeed, the question of the tightness of the rightmost leaf in the extremely heavy-tailed case remains an open problem.

### v2

@@ 4.1 Moderate tails
Since $Y_{n}\to\infty$ in probability, $I_{n}\to\infty$ in probability as well. Moreover, for all $n\geq 1$, one can choose $N_{n}$ such that $N_{n}-n$ is distributed as $N_{0}$. The fact that all vertices $k\geq N_{n}$ belong to the tree rooted at $I_{n}$ in this single-choice model implies that $I_{n}\in T_{k}$ for all $k\geq N_{n}$, and in particular $_{I_{n}}\leq{}_{k}$ for all $k\geq N_{n}$. From Proposition 11 we know that $_{n}\to\infty$ in probability. Hence, $_{I_{n}}\to\infty$ in probability. Moreover, for any fixed $m\geq 1$ we have
$\mathbb{P}\left(\cup_{k\geq 2n}\left[{}_{k}\leq m\right]\right)$ $\leq\mathbb{P}\left(N_{n}>2n\right)+\mathbb{P}\left(N_{n}\leq 2n,{}_{I_{n}}\leq m\right)$
$\leq\mathbb{P}\left(N_{0}>n\right)+\mathbb{P}\left({}_{I_{n}}\leq m\right)~.$
Thus, for any $m\geq 1$, $\mathbb{P}\left(\cup_{k\geq 2n}[{}_{k}\leq m]\right)\to 0$ as $n\to\infty$. By continuity of measure, this implies that for any $m\geq 1$,
$\mathbb{P}\left(\cap_{n\geq 1}\cup_{k\geq 2n}\left[{}_{k}\leq m\right]\right)=0.$
@@
$\mathbb{P}\left(\cup_{m\geq 1}\cap_{n\geq 1}\cup_{k\geq 2n}\left[{}_{k}\leq m\right]\right)=0~,$
which implies that $_{n}\to\infty$ almost surely.
@@ 4.2 Heavy tails
In this section, we study the heavy-tailed case, that is, $\mathbb{E}\min(Z,W)=\infty$. From Theorem 3 we already know that n cannot go to infinity with probability one. Here we show that, in contrast to this, $_{n}\to\infty$ in probability when $\sumop\displaylimits_{n=1}^{\infty}r_{n}^{2}<\infty$.
@@
$\mathbb{P}\left(-k\in T_{n}^{Z}\right)$ $=\sumop\displaylimits_{i=1}^{n}\mathbb{P}\left(Z_{i}=i+k\right)\mathbb{P}\left(i\in T_{n}\right)$
$=\sumop\displaylimits_{i=1}^{n}q_{k+i}r_{n-i}$
$=\sumop\displaylimits_{i=k+1}^{n+k}q_{i}r_{n+k-i}$
$\leq\sumop\displaylimits_{i=1}^{n+k}q_{i}r_{n+k-i}=r_{n+k}~.$
Proof. [Proof of Theorem 7] Assume that $\sumop\displaylimits_{n=1}^{\infty}r_{n}^{2}<\infty$. Fix a constant $w>0$ and let $I_{n}^{w}$ denote the number of intervals of the form $(kw,(k+1)w]$, with $k\in\mathbb{Z}_{\geq 0}$, that are intersected by the chain $T_{n}^{Z}$. Note that, for any $x>0$, if $I_{n}^{w}<x$ then $T_{n}^{Z}<wx$, thus
@@
$\quad\leq\mathbb{E}\sumop\displaylimits_{\begin{subarray}{c}k,m\in A_{n}\\ \textit{s.t. }k\neq m\end{subarray}}\mathbbm{1}_{T_{k}^{W}\cap T_{m}^{W}\neq\emptyset}$
$\quad\leq\mathbb{E}\sumop\displaylimits_{\begin{subarray}{c}k,m\in A_{n}\\ \textit{s.t. }k\neq m\end{subarray}}\mathbb{P}\left(T_{k}^{W}\cap T_{m}^{W}\neq\emptyset\right)$
@@
$\mathbb{P}\left(T_{k}^{W}\cap T_{m}^{W}\neq\emptyset\right)$ $\leq\sumop\displaylimits_{i=-\infty}^{k}\mathbb{P}\left(i\in T_{k}^{W}\right)\mathbb{P}\left(i\in T_{m}^{Z}\right)$
$\leq\sumop\displaylimits_{i=-\infty}^{k}r_{k-i}r_{m-i}$
$\leq\sqrt{\sumop\displaylimits_{i=0}^{\infty}r_{i}^{2}}\times\sqrt{\sumop\displaylimits_{i=m-k}^{\infty}r_{i}^{2}}.$ (4.19)
@@
$\mathbb{P}\left(\cup_{k\neq m,k\in A_{n}}T_{k}^{W}\cap T_{m}^{W}\neq\emptyset\right)\leq{\ell\choose 2}\sqrt{\sumop\displaylimits_{i=0}^{\infty}r_{i}^{2}}\times\sqrt{\sumop\displaylimits_{i=w}^{\infty}r_{i}^{2}}.$ (4.20)
@@
Finally, observe that if $_{n}$$\ell$ and $I_{n}^{w}\geq 2\ell$, then there have to be at least two distinct vertices $k,m$ in $A_{n}$ such that the chains $T_{k}^{W}$ and $T_{m}^{W}$ meet each other: otherwise, each chain $T_{k}^{W}$ with $k\in A_{n}$ would lead to a distinct leaf in $T_{n}$, contradicting the event $_{n}$$\ell$. Thus, by (4.20) we have
$\mathbb{P}\left({}_{n}<\ell\right)\leq\mathbb{P}\left(I_{n}^{w}<2\ell\right)+\epsilon~.$
@@
$\limsup_{n\to\infty}\mathbb{P}\left({}_{n}<\ell\right)\leq\epsilon~,$
@@
5 Generalization to $k$ choices.
A natural generalization of the model allows for $k\geq 2$ choices. Fix $k\geq 2$ and consider $k$ independent sequences of i.i.d. random variables $(Z_{i}^{(1)})_{i\geq 1},\dots,(Z_{i}^{(k)})_{i\geq 1}$. Given a sequence of preference values $(U_{i})_{i\leq 0}$, we define a sequential random coloring by setting $V_{n}=U_{n}$ and $C_{n}=n$ for all $n\leq 0$, and for $n>0$ letting
$C_{n}=C_{n-Z_{n}^{(j)}}\quad\text{and}\quad V_{n}=V_{n-Z_{n}^{(j)}}\quad\text{where}\quad j=\mathop{\mathrm{argmin}}_{1\leq i\leq k}V_{n-Z_{n}^{(i)}}$
for $n\geq 1~.$ The set of vertices reached from vertex $n$ is now defined recursively as
$T_{n}=\{n\}\cup\left(\bigcupop\displaylimits_{i=1}^{k}T_{n-Z_{n}^{(i)}}\right),\quad n\geq 1~,$
with $T_{n}=\{n\}$ for all $n\leq 0$. As in the two-choice model, we may then define $R_{n}$, n, and $L_{n}$ accordingly. The graph associated with this model therefore, contains multiple embedded copies of the two-choice process. It follows from this observation that consistency in the two-choice model carries over to the $k$-choice model for any $k\geq 2$, and for each of the three configurations of the initial pool of topics considered in this paper. As a consequence, Theorem 4 extends automatically to the $k$-choice model:
Theorem 15.
Assume that $q_{i}>0$ for all $i\in\mathbb{N}$ and $\mathbb{E}Z=\infty$. Then, in the $k$-choice model, with probability one,
$L_{n}\to-\infty~.$
Thus, the system is strongly consistent in configuration (ii).
However, the increase in the number of available recommendations allows one to work under weaker assumptions on the tail of $Z$, and in this setting the results can be sharpened. In particular, one can adapt the proof given in Section 4.1 by modifying the recursion defining $S_{n}$ to
$S_{n}=\{n\}\cup S_{\,n-\min_{1\leq i\leq k}Z_{n}^{(i)}},$
and by controlling the growth of
$J_{n}=\sumop\displaylimits_{m=1}^{n}\mathbbm{1}_{m\in S_{n}}\,\mathbbm{1}_{\max_{1\leq i\leq k}Z_{m}^{(i)}\geq m}~,$
to establish the following result.
Theorem 16.
Assume that the distribution of $Z$ is aperiodic, and that
$\mathbb{E}[Z]=\infty,\qquad\mathbb{E}\!\left[\min_{1\leq i\leq k}Z^{(i)}\right]<\infty~,$
where $Z^{(1)},\dots,Z^{(k)}$ are independent copies of $Z$. Then, in the $k$-choice model,
$_{n}\to\infty\quad\text{almost surely.}$
The arguments used in the proofs of Theorems 2, 3 and Theorem 5 extend naturally to the present setting, leading to the following three results.
Theorem 17.
Let $L_{\infty}=\inf_{n\in\mathbb{N}}L_{n}$. If $\mathbb{E}Z<\infty$, then $L_{\infty}$ is finite almost surely. Thus, in the $k$-choice model, the sequences $(R_{n})_{n\in\mathbb{N}},(L_{n})_{n\in\mathbb{N}}$ and $({}_{n})_{n\in\mathbb{N}}$ are almost surely bounded.
Theorem 18.
Let $Z^{(1)},\dots,Z^{(k)}$ be independent copies of $Z$. Assume that
$\mathbb{E}\min_{1\leq i\leq k}Z^{(i)}=\infty~.$
Then, in the $k$-choice model, with probability one,
$\liminf_{n\to\infty}{}_{n}$$\infty~$.
Theorem 19.
Let $Z^{(1)},\dots,Z^{(k)}$ be independent copies of $Z$.
  1. 1.
    If $\mathbb{E}\min_{1\leq i\leq k}Z^{(i)}<\infty$, then $\sup_{n\in\mathbb{N}}|R_{n}|<\infty$ almost surely in the $k$-choice model.
  2. 2.
    If $\mathbb{E}\min_{1\leq i\leq k}Z^{(i)}=\infty$, then $\sup_{n\in\mathbb{N}}|R_{n}|=\infty$ almost surely in the $k$-choice model.
In summary, the consistency table from Section 1.4 remains unchanged in the $k$-choice model. The only difference is that the moderate-tail regime is now characterized by the conditions $\mathbb{E}\min_{1\leq i\leq k}Z^{(i)}<\infty$ and $\mathbb{E}Z^{(1)}=\infty$, while the heavy-tail regime corresponds to $\mathbb{E}\min_{1\leq i\leq k}Z^{(i)}=\infty$.
@@ 6 Conclusion
We introduced and studied a simple mathematical model for recommendation systems based on giving each user two random choices, which include a mixture of past choices and untried options. In this setup, we identify three regimes based on the tail behavior of the sizes of the jumps in the past. Although our model is primarily theoretical, its analysis offers qualitative insights that may inform the design of recommendation systems. In particular, it highlights how structural properties of the recommendation process can influence long-term behavior and diversity. Extending the model to incorporate personalization and adaptive feedback remains an interesting direction for future work.
6.1 Comparison with standard baselines
To situate our results within the broader landscape of recommendation methods, it is natural to consider connections to standard algorithmic baselines. Classical approaches, such as collaborative filtering, multi-armed bandits, and deep learning–based recommender systems, rely on adaptive strategies to cope with noisy or partially observed user preferences. While our model is highly stylized and focuses on the structural properties of the recommendation process rather than optimization or learning, comparing its behavior conceptually to these frameworks helps highlight both its scope and its limitations.
A natural point of comparison is with multi-armed bandit algorithms [16, 17], which also model sequential decision-making and recommendation dynamics. However, the two frameworks differ fundamentally: in our model, the “rewards” (preference values) are static and noise-free, with randomness arising solely from the distribution of edge lengths, while in bandit algorithms, the primary challenge is to manage uncertainty in the reward distributions through exploration and exploitation. Because of this distinction, standard bandit baselines are not directly applicable to our setting, and a meaningful comparison would require a reformulation of the bandit problem that departs significantly from its usual assumptions.
A closer comparison can be drawn with bandit algorithms in the noiseless setting, where the reward associated with each arm is fixed and known after a single pull. In that framework, the problem is essentially trivial: the optimal arm is identified immediately, and the learning phase vanishes. By contrast, our model remains nontrivial even in the absence of noise, since the randomness lies not in the evaluation of topics but in the structure of the recommendation process itself. The stochastic nature of the edge lengths forces the system to repeatedly revisit or discover new topics, giving rise to rich asymptotic behaviors such as strong or weak consistency.
6.2 Perspectives and open problems
One might naturally expect the $k$-choice model to exhibit distinct behavior in the $k+1$ regimes defined by the tail conditions
$\mathbb{E}\min_{1\leq i\leq\ell}Z^{(i)}=\infty\quad\text{and}\quad\mathbb{E}\min_{1\leq i\leq\ell+1}Z^{(i)}<\infty,$
for each $\ell\in\{0,\dots,k\}$ (with the understanding that the left-hand condition is omitted when $\ell=0$ and the right-hand condition is omitted when $\ell=k$). Although our results extend straightforwardly to the $k$-choice setting, the current consistency analysis does not distinguish between these individual regimes. A deeper understanding of the distinctions between these regimes and their implications for recommendation outcomes would be of significant interest.
The present model remains highly abstract and does not directly reflect real-world applications. For instance, the assumption of static preference values for each topic may not capture the evolution of actual user behavior. A natural direction for future work is to develop dynamic generalizations of the model that account for time-varying preferences for each topic.
Appendix
Symbol Description
$T_{n}$ Set of vertices reached by vertex $n$.
$\mathcal{L}_{n}$ Set of vertices with non-positive index (or leaves) that is reached from vertex $n$. ($\mathcal{L}_{n}=T_{n}\cap\{0,-1,-2,\dots\})$
n Number of leaves reached from $n$. ($_{n}=|\mathcal{L}_{n}|$)
$R_{n}$ Rightmost leaf in $T_{n}$. $(R_{n}=\max\left\{i:i\in\mathcal{L}_{n}\right\})$
$L_{n}$ Leftmost leaf in $T_{n}$. ($L_{n}=\min\left\{i:i\in\mathcal{L}_{n}\right\}$)
$T_{n}^{Z}$ (resp. $T_{n}^{W})$ Chain in $T_{n}$ obtained by repeatedly taking the $Z$-edges (resp. the $W$-edges)
$S_{n}$ Chain in $T_{n}$ obtained by repeatedly taking shortest outgoing edges
$J_{n}$ Number of vertices in $S_{n}$ connected to a leaf via their longest outgoing edge
$C_{n}$ Topic recommended at time $n$
$Z_{n},W_{n}$ Independent random delays for recommendations
$U_{i}$ Preference value of topic $i$ ($i\leq 0$)
$V_{n}$ Preference value of topic seen at time $n$
$p_{i},q_{i}$ Probabilities associated with $Z$ ($p_{i}=\mathbb{P}(Z\geq i)$, $q_{i}=\mathbb{P}(Z=i)$)
$p_{i}^{*}$ Upper tail probability associated with $\max(Z,W)$ ($p_{i}^{*}=\mathbb{P}(\max(Z,W)\geq i)=2p_{i}-p_{i}^{2}$)
$r_{n}$ Probability that vertex 0 belongs to the $Z$-chain starting at vertex $n$ ($r_{n}=\mathbb{P}\left(0\in T_{n}^{Z}\right)=\mathbb{P}\left(0\in T_{n}^{W}\right))$
$v_{n}$ Probability that vertex 0 belongs to the chain $S_{n}$ ($v_{n}=\mathbb{P}\left(0\in S_{n}\right)$)
$\mathbbm{1}_{A}$ Indicator function of event AA
$\mathbb{E}X$ Expected value of random variable $X$
$\mathbb{V}X$ Variance of random variable $X$
Notation Table
Funding
Luc Devroye was supported by NSERC under grant number RGPIN-2024-04164. Gábor Lugosi acknowledges the support of Ayudas Fundación BBVA a Proyectos de Investigación Científica 2021 and the Spanish Ministry of Economy and Competitiveness grant PID2022-138268NB-I00, financed by MCIN AEI 10.13039 501100011033, FSE+MTM2015-67304-P, and FEDER, EU.